Let X represent the full height of a certain species of tree. Assume that X has a normal probability distribution with mean 219.5 ft and standard deviation 62.5 ft You intend to measure a random sample of n 187 trees. The bell curve below represents the distibution of these sample means. The scale on the horizontal axis is the standard error of the sampling distribution Complete the indicated boxes, correct to two decimal places Points possible: 1 Unlimited attempts 8:11 PM 1/22/2019 EWLETTPACKARD BANG OLOFSE prt sc 5 6 8 9 +back Please help ASAP please.
In a recent year, the Better Business Bureau settled 75% of complaints they received. (Source: USA Today, March 2 2009) You have been hired by the Bureau to investigate complaints this year involving computer stores. You plan to select a random sample of complaints to estimate the proportion of complaints the Bureau is able to settle. Assume the population proportion of complaints settled for the computer stores is the 0.75, as mentioned above. Suppose your sample size is 159. What is the probability that the sample proportion will be within 3 percent of the population proportion? Note: You should carefully round any z-values you calculate to 4 decimal places to match wamaps approach and calculations. Answer Enter your answer as a number accurate to 4 decimal places.) Points possible: 1 Unlimited attempts Subemit 8:12 1/22/ HEWLETT PACKARD prt sc 6 7 8 9 O P
A Food Marketing Institute found that 25% of households spend more than S125 a week on groceries. Assume the population proportion is 0.25 and a simple random sample of 211 households is selected from the population. What is the probability that the sample proportion of households spending more than $125 a week is less than 0.272 Note: You should carefully round any z-values you calculate to 4 decimal places to match wamaps approach and Answer (Enter your answer as a number accurate to 4 decimal places.) Points possible: 1 812 HEWLETTPACKARD ort sc 6 8 9 T Y U
Nizam Hami A Food Marketing Institute found that 56% of households spend more than $125 a week on groceries. Assume the population proportion is 0.56 and a simple random sample of 127 households is selected from the population. What is the probability that the sample proportion of households spending more than $125 a week is more than than 0.36? Note: You should carefully round any z-values you calculate to 4 decimal places to match wamaps approach and calculations. Enter your answer as a number accurate to 4 decimal places.) Answer- Points possible: 1 Unlimited attempts. t icense Submit BANG l44 prt sc 6 8 9 0
A leading magazine (like Barrons) reported at one time that the average number of weeks an individual is unemployed is 30 weeks. Assume that for the population of all unemployed individuals the population mean length of unemployment is 30 weeks and that the population standard deviation is 4 weeks. Suppose you would like to select a random sample of 34 unemployed individuals for a follow-up study Find the probability that a single randomly selected value is less than 3 nr-31)- Find the probability that a sample of size n- 34 is randomly selected with a mean less than 31. PM< 31)- Enter your answers as numbers accurate to 4 decimal places. Points possible: 1 Liceme Unlimited attempts. Submit EWLETT PACKARD ANG &O prt sc 6 8 9 O P
A leading magazine (like Barrons) reported at one time that the average number of weeks an individual is unemployed is 39 weeks. Assume that for the population of all unemployed individuals the population mean length of unemploymen is 39 weeks and that the population standard deviation is 2.8 weeks. Suppose you would like to select a random sample of 85 unemployed individuals for a follow-up study Find the probability that a single randomly selected value is greater than 38.4 PLX> 384) (Enter your answers as numbers accurate to 4 decimal places.) Find the probability that a sample of size n 85 is randomly selected with a mean greater than 38.4 PM> 38.4) Enter your answers as numbers accurate to 4 decimal places.) Points possible: 1 License Submit ARD BANG &О 10 ort sc 6 7 8 9 0 Y U O P
A leading magazine (like Barrons) reported at one time that the average number of weeks an individual is unemployed is 13.2 weeks. Assume that for the population of all unemployed individuals the population mean length of unemployment is 13.2 weeks and that the population standard deviation is 8.8 weeks. Suppose you would like to select a random sample of 29 unemployed individuals for a follow-up study Find the probability that a single randomly selected value is between 8 and 14.2 PI8<X 14.2) Find the probability that a sample of size n 29 is randomly selected with a mean between 8 and 142. Enter your answers as numbers accurate to 4 decimal places. Liceme Points possible: 1 Unlimited attempts. 1/22 10 ort sc 6 8 9 T YU
Answer #1

Solution :- Given that mean = 219.5, standard deviation = 62.5 , n = 187

μx = 219.5 σx = 62.5/sqrt(187) = 4.57

=> Lower limit : 219.5 - (2*4.57) = 210.36

=> middle : 219.5

=> upper limit : 219.5 + (2*4.57) = 228.64

-----------------

Solution:-

That is we need sample proportion between 0.75-0.03 = 0.72 and 0.75+0.03 = 0.78.
Here we have p=0.75 and n=159. So z-score for sample proporiton p^ =0.72 is

Z = (p^-p)/sqrt((pq/n)) = (0.72-0.75)/sqrt(0.75*0.25/159) = -0.8736

z-score for sample proporiton p^ = 0.78

Z = (p^-p)/sqrt((pq/n)) = (0.78-0.75)/sqrt(0.75*0.25/159) = 0.8736

=> P(-0.8736 < Z < 0.8736) = 0.6156

-------------------------